40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jun 26 21:01:26 2019, gautam, Update, LSC, POX and POY locking 
    Reply  Thu Jun 27 14:28:12 2019, gautam, Update, LSC, POX and POY locking 
       Reply  Tue Jul 2 17:34:54 2019, gautam, Update, LSC, Irides opened up on EY table 
          Reply  Tue Jul 9 18:06:05 2019, gautam, Update, LSC, Y-arm ASS in a workable state ditherFreqs.pdftransRenorm.png
Message ID: 14704     Entry time: Wed Jun 26 21:01:26 2019     Reply to this: 14705
Author: gautam 
Type: Update 
Category: LSC 
Subject: POX and POY locking 

Now that the IMC is remaining locked for extended periods of time, the next problem to attack is the ASS dither alignment system. For a start, I decided to try and get the POX and POY locking working again, as we have not fully recovered the interferometer alignment after the most recent pumpdown. I spent a couple of hours tweaking the alignment of the arm cavity mirrors, BS, and TTs to try and recover the maximum possible TRX and TRY - however, my best efforts only yielded TRX~0.8, TRY~0.75. Moreover, the beam axis is such that the spot is significantly off in YAW on both ETMs, as evidenced by the camera views (also true but less obvious on the ITMs). However, trying to bring the beam back to the center of the optics yields TRY and TRX values lower than the above reported maxima. The EX green beam is currently unavailable to verify the arm cavity alignment because of my hijacking the EX NPROs PZT control for PLL investigations, but with the Y arm, I'm able to lock a TEM00 mode. Probably just needs more careful systematic alignment, but I'm not pursuing this tonight.

ELOG V3.1.3-